OpenStudy (anonymous):

Does the function f(x)= x^2-2x+2 have a maximum or minimum value?

OpenStudy (raden):

see the coefficient of x^2 is positive, so it has minimum value but if the coefficient of x^2 is negative then it has maximum value

OpenStudy (anonymous):

No the x^2 is not negative the x^2 means the 2 is the exponent . So it's suppose to read x squared.

OpenStudy (raden):

the general equation of function quadratic is f(x) = ax^2 + bx + c see the equation above is f(x) = x^2 - 2x + 2, here known a =1 , b = -2, and c = 2. so a is positive, right ?